# Where Do Passengers Gaze? Impact of Passengers’ Personality Traits on Their Gaze Pattern Toward Pedestrians During APMV-Pedestrian Interactions with Diverse eHMIs

### v1

@@ Abstract
Autonomous Personal Mobility Vehicles (APMVs) are designed to address the “last-mile” transportation challenge for everyone. When an APMV encounters a pedestrian, it uses an external Human-Machine Interface (eHMI) to negotiate road rights. Through this interaction, passengers also engage with the process. This study examines passengers’ gaze behavior toward pedestrians during such interactions, focusing on whether different eHMI designs influence gaze patterns based on passengers’ personality traits. The results indicated that when using a visual-based eHMI, passengers often struggled to perceive the communication content. Consequently, passengers with higher Neuroticism scores, who were more sensitive to communication details, might seek cues from pedestrians’ reactions. In addition, a multimodal eHMI (visual and voice) using neutral voice did not significantly affect the gaze behavior of passengers toward pedestrians, regardless of personality traits. In contrast, a multimodal eHMI using affective voice encouraged passengers with high Openness to Experience scores to focus on pedestrians’ heads. In summary, this study revealed how different eHMI designs influence passengers’ gaze behavior and highlighted the effects of personality traits on their gaze patterns toward pedestrians, providing new insights for personalized eHMI designs.
@@ I-A Background
Autonomous personal mobility vehicle (APMV) is a small autonomous vehicle designed to solve the last-mile problem for everyone (not only for the elderly or people with disabilities). APMV is equipped with SAE Level 3 to 5 automated driving systems and can drive autonomously in shared spaces (see Fig. 1). In such driving scenarios, APMVs frequently encounter pedestrians and interact with them, even negotiating the right-of-way [1, 2]. To improve communication between APMVs and pedestrians, external human-machine interface (eHMI), which is widely used in pedestrian and autonomous vehicle (AV) research [3, 4], is also applied to APMVs [5, 1, 6].
@@ I-B Challenges in eHMI Design for APMV Passengers
Different from cars, APMVs often have an open-body design. Therefore, when an APMV uses eHMI to communicate with pedestrians, its passenger is also exposed to these interactions. Based on Edward T. Hall’s Proxemics Theory [7], in pedestrian-APMV interactions, as shown in Fig. 1, the distance between passengers and pedestrians is within the social distance range (1.2-3.6 meters). According to [7], this range gives enough personal space and allows for clear communication between the passenger and the pedestrian APMV. For this unique interaction among the pedestrian, the APMV, and the APMV passenger, it is important to consider both the efficiency of communication with pedestrians and the user experience of passengers when designing eHMI for APMVs [1].
For this challenge, Zhang et al. [6] suggested that when APMVs interact with pedestrians, the APMV should share its driving intention with both pedestrians and its passenger via an eHMI. Thus, they proposed using a projector-type eHMI to display the driving path on the ground to show its driving intentions. In our previous study, Liu et al. [1] suggested a multimodal eHMI for APMV that uses LED lights to project colors under the APMV to show its driving intentions, and designed facial expressions and voices to communicate with pedestrians through a display and a speaker. Moreover, Liu et al. [1] analyzed subjective feelings of passengers during interactions between APMV and pedestrians. They found that passengers with different personality traits have different preferences for voice in eHMIs and recommended considering these personality differences to design APMV eHMIs. However, there is still a lack of clear guidelines on how to customize the eHMI design based on the personality of the passengers.
@@ II RELATED WORKS
In the transportation field, gaze behavior has been widely studied in pedestrian crossing scenarios. For example, by analyzing pedestrians’ gaze behavior to understand their crossing strategies, researchers have explored how pedestrians allocate visual attention during road-crossing tasks and how gaze behavior influences decision-making [11]. Furthermore, study [12] had examined how normally sighted pedestrians use gaze to navigate and safely cross various types of intersection.
@@ III-A Participants
24 Japanese participants (12 males and 12 females, self-reported) participated in the ride experiment as APMV passengers. Each participant took part in the experiment for one hour and received 1,000 JPY as a reward.
@@ III-B Autonomous Personal Mobility Vehicle (APMV)
As shown in Fig. 1, an APMV is developed based on the WHILL Model CR, which is equipped with a LiDAR (Velodyne VLP-16) and a control computer. An autonomous driving system with SAE Level 3 was developed based on ROS Noetic iiiROS Noetic: http://wiki.ros.org/noetic/. This system uses LiDAR to detect obstacles, avoid them or stop, and automatically follow a predefined path.
In this experiment, the maximum speed was set to 1 m/s (3.6 km/h) for safety reasons. Moreover, the APMV’s stopping and starting during pedestrian interactions were controlled by an experimenter behind the scenes.
@@ III-E2 Passenger’s gaze duration on pedestrian’s head
As shown in Fig. 4, an Area of Interest (AOI) is defined to analyze passengers’ gaze behavior toward the pedestrians’ head. More specifically, the DensePose labels “HEAD_LEFT” and “HEAD_RIGHT” used to define the AOI are shown in Fig. 4. The red circle, with a diameter of 50 pixels, represents the gaze position. The areas covered by the red circle that overlap with the AOI regions are classified as gazes on it. As shown in Fig. 4, if the DensePose prediction output contains multiple prediction results, any result that includes either “HEAD_LEFT” and “HEAD_RIGHT” is defined as the gaze on the head.
@@ IV-B Gaze duration under eHMI conditions
IV-C Partial Correlations Analysis
Considering that the five traits of the passengers’ personality jointly influence their gaze behavior, we used partial correlation analysis to explore the relationship between each personality traits and gaze behavior under different eHMI conditions.
Tables III, IV and V show the results of the partial correlation analysis between passengers’ personality traits and gaze durations on pedestrians’ head during APMV interactions with pedestrians using the eHMI-T, eHMI-NV, and eHMI-AV, respectively.
The above results show that when the APMV uses the eHMI-T, there is a significant weak positive partial correlation between participants’ Neuroticism scores and their gaze duration on head ($r=0.206,p=0.047$).
When using eHMI-NV, no significant partial correlation was found between the Big Five personality traits of the passengers and their gaze duration on pedestrians’ head.
When the APMV uses the eHMI-AV, a significant weak positive partial correlation was found between passengers’ Openness to Experience scores and their gaze duration on pedestrians’ head ($r=0.220,p=0.032$).
TABLE III: Partial correlations between passengers’ personality traits and their gaze duration on pedestrians’ head during APMV interactions with pedestrians using the eHMI-T.
@@ X Y n r CI 95% p BF10 power
n: number of trials. *:$p<0.05$.
: Gaze data from 3 out of 4 trials for one participant were defective, so these data of 3 trials were excluded.
TABLE IV: Partial correlations between passengers’ personality traits and their gaze duration on pedestrians’ head during APMV interactions with pedestrians using the eHMI-NV.
@@
TABLE V: Partial correlations between passengers’ personality traits and their gaze duration on pedestrians’ head during APMV interactions with pedestrians using the eHMI-AV.
@@ V-A Effects of eHMI-T on passenger gaze behaviors
From the partial correlation analysis results shown in Table III, when the APMV communicates with pedestrians using eHMI-T, there was only a significant partial correlation relationship was observed between the passengers’ Neuroticism scores and their gaze duration on the pedestrians’ head with a positive partial correlation. This result suggests that passengers with lower Neuroticism tended to avert their gaze from the pedestrian’s head while immersed in the interaction with the pedestrian and the silent eHMI-T, whereas passengers with higher Neuroticism tended to allocate longer visual attention to the pedestrian’s head.
However, studies [22, 23] suggested that people with a high Neuroticism score tended to look away when co-communicators looked them in the eyes during a conversation or other social situations. We considered that this difference arises because the studies [22, 23] focused on face-to-face interactions, where mutual eye contact was more likely to occur. In our experiment, the pedestrian primarily interacted with the eHMI and did not look at the passenger throughout the process. Moreover, since passengers could not directly access the communication content from the eHMI-T when the APMV interacted with pedestrians, they had to infer it based on the pedestrians’ reactions. Given this uncertainty, we considered that passengers with higher neuroticism scores might remain more vigilant in such an ambiguous communication environment. Studies [24, 25] provided some thread showing that people with anxious personalities, such as high Neuroticism, exhibit increased vigilance toward potentially threatening stimuli.
@@ V-B Effects of eHMI-NV on passenger gaze behaviors
From the results of the partial correlation analysis shown in Table IV, when using eHMI-NV, no significant correlation was found between passengers’ Big Five personality traits and their gaze behavior. This result suggested that the neutral communication style of eHMI-NV provided a relatively consistent experience for passengers with different personality traits, and it was difficult to affect the gaze behavior of passengers with specific personality traits.
@@ V-D Limitations
The participants in this study are Japanese, and cultural factors may have influenced their specific gaze behavior. The results may be limited by the specific eHMI designs used in this study. The pedestrians in the experiment were played by experimenters. Their behaviors and reactions in the interaction scenarios were not as varied as those of real pedestrians. This limitation in diversity may potentially affect the results of this study.
@@ VI CONCLUSION
This study investigated the differences in gazing behavior towards pedestrians’ head by APMV passengers with different personalities when APMV communicates with pedestrians using three different types of eHMI The results indicated that when using the eHMI-T, passengers often struggled to perceive the communication content. Consequently, passengers with higher Neuroticism scores, who were more sensitive to communication details, might seek cues from pedestrians’ reactions. In addition, eHMI-NV using neutral voice did not significantly affect the gaze behavior of passengers toward pedestrians, regardless of personality traits. In contrast, eHMI-AV using affective voice encouraged passengers with high Openness to Experience scores to focus on pedestrians’ heads.

### v2

@@ Abstract
Autonomous Personal Mobility Vehicles (APMVs) are designed to address the “last-mile” transportation challenge for everyone. When an APMV encounters a pedestrian, it uses an external Human-Machine Interface (eHMI) to negotiate road rights. Through this interaction, passengers are also passively exposed to the process. This study examines passengers’ gaze behavior toward pedestrians during such interactions, focusing on whether passengers’ personality traits influence their gaze patterns towards pedestrians when using different eHMI designs. When using a visual-based eHMI, which caused passengers to struggle in perceiving the communication content, the results suggested that passengers with higher Neuroticism scores, who were more sensitive to communication details, might seek cues from pedestrians’ reactions. In addition, a multimodal eHMI (visual and voice) using neutral voice did not significantly affect the gaze behavior of passengers toward pedestrians, regardless of personality traits. In contrast, a multimodal eHMI using affective voice encouraged passengers with high Openness to Experience scores to focus on pedestrians’ heads. In summary, this study revealed how different eHMI designs influence passengers’ gaze behavior and highlighted the effects of personality traits on their gaze patterns toward pedestrians, providing new insights for personalized eHMI designs.
@@ I-A Background
Autonomous personal mobility vehicle (APMV) is a small autonomous vehicle designed to solve the last-mile problem for everyone (not only for the elderly or people with disabilities). APMV is equipped with SAE Level 3 to 5 automated driving systems and can drive autonomously in shared spaces (see Fig. 1). In such driving scenarios, APMVs frequently encounter pedestrians and interact with them, even negotiating the right-of-way [1, 2]. To improve communication between APMVs and pedestrians, external human-machine interface (eHMI), which is widely used in pedestrian and autonomous vehicle (AV) research [3, 4, 5], is also applied to APMVs [6, 1, 7].
@@ I-B Challenges in eHMI Design for APMV Passengers
Different from cars, APMVs often have an open-body design. Therefore, when an APMV uses eHMI to communicate with pedestrians, its passenger is also passively exposed to these interactions. Based on Edward T. Hall’s Proxemics Theory [8], in pedestrian-APMV interactions, as shown in Fig. 1, the distance between passengers and pedestrians is within the social distance range (1.2-3.6 meters). According to [8], this range gives enough personal space and allows for clear communication between the passenger and the pedestrian APMV. For this unique interaction among the pedestrian, the APMV, and the APMV passenger, it is important to consider both the efficiency of communication with pedestrians and the user experience of passengers when designing eHMI for APMVs [1].
For this challenge, Zhang et al. [7] suggested that when APMVs interact with pedestrians, the APMV should share its driving intention with both pedestrians and its passenger via an eHMI. Thus, they proposed using a projector-type eHMI to display the driving path on the ground to show its driving intentions. In our previous study, Liu et al. [1] suggested a multimodal eHMI for APMV that uses LED lights to project colors under the APMV to show its driving intentions, and designed facial expressions and voices to communicate with pedestrians through a display and a speaker. Moreover, Liu et al. [1] analyzed subjective feelings of passengers during interactions between APMV and pedestrians. They found that passengers with different personality traits have different preferences for voice in eHMIs. Thus, they recommended considering passengers’ personality differences to design eHMI of APMVs. However, there is still a lack of clear guidelines on how to customize the eHMI design based on the personality of the passengers.
@@ II RELATED WORKS
In the transportation field, gaze behavior has been widely studied in pedestrian crossing scenarios. For example, by analyzing pedestrians’ gaze behavior to understand their crossing strategies, researchers have explored how pedestrians allocate visual attention during road-crossing tasks and how gaze behavior influences decision-making [12]. Furthermore, study [13] had examined the gaze behavior of normally sighted pedestrians when crossing various types of intersection safely.
@@ III-A Participants
24 Japanese participants (12 males and 12 females, self-reported; 22 to 29 years with mean 23.9 years and standard deviation 1.7 years) participated in the ride experiment as APMV passengers. All participants had no experience with APMV as well as eHMI. Each participant took part in the experiment for one hour and received 1,000 JPY as a reward.
@@ III-B Autonomous Personal Mobility Vehicle (APMV)
As shown in Fig. 1, an APMV is developed based on the WHILL Model CR, which is equipped with a LiDAR (Velodyne VLP-16) and a control computer. An autonomous driving system (ADS) with SAE Level 3 was developed based on ROS Noetic iiiROS Noetic: http://wiki.ros.org/noetic/. This system uses LiDAR to detect obstacles, avoid them or stop, and automatically follow a predefined path.
Even though the APMV has a Level 3 ADS and can autonomously drive on a pre-planned route, for safety reasons, the APMV’s stopping and starting during pedestrian interactions were controlled by an experimenter behind the scenes. Furthermore, the maximum speed was limited to 1 m/s (3.6 km/h) for the same safety reasons.
@@ III-E2 Passenger’s gaze duration on pedestrian’s head
As shown in Fig. 4, an Area of Interest (AOI) is defined to analyze passengers’ gaze behavior toward the pedestrians’ head. More specifically, the DensePose labels “HEAD_LEFT” and “HEAD_RIGHT” used to define the AOI are shown in Fig. 4. The red circle, with a diameter of 50 pixels, represents the gaze position. The areas covered by the red circle that overlap with the AOI regions are classified as gazes on it. As shown in Fig. 4, if the DensePose prediction output contains multiple prediction results, any result that includes either “HEAD_LEFT” and “HEAD_RIGHT” is also defined as the gaze on the head.
@@ IV-B Gaze duration under eHMI conditions
IV-C Pairwise Pearson Correlations Analysis
Pairwise pearson correlation analysis to explore the linear relationship between each personality traits and gaze behavior under different eHMI conditions.
Tables III, IV and V show the results of the pairwise correlation analysis between passengers’ personality traits and gaze durations on pedestrians’ head during APMV interactions with pedestrians using the eHMI-T, eHMI-NV, and eHMI-AV, respectively.
The above results show that when the APMV uses the eHMI-T, there is a significant weak positive correlation between participants’ Neuroticism scores and their gaze duration on head ($r=0.206,p=0.047$).
When using eHMI-NV, no significant correlation was found between the Big Five personality traits of the passengers and their gaze duration on pedestrians’ head.
When the APMV uses the eHMI-AV, a significant weak positive correlation was found between passengers’ Openness to Experience scores and their gaze duration on pedestrians’ head ($r=0.220,p=0.032$).
TABLE III: Pearson correlations between passengers’ personality traits and their gaze duration on pedestrians’ head during APMV interactions with pedestrians using the eHMI-T.
@@ X Y n r CI 95% p BF10 power
n: number of trials. *:$p<0.05$.     : Gaze data from 3 out of 4 trials for one participant were defective, so these data of 3 trials were excluded.
TABLE IV: Pearson correlations between passengers’ personality traits and their gaze duration on pedestrians’ head during APMV interactions with pedestrians using the eHMI-NV.
@@
TABLE V: Pearson correlations between passengers’ personality traits and their gaze duration on pedestrians’ head during APMV interactions with pedestrians using the eHMI-AV.
@@ V-A Effects of eHMI-T on passenger gaze behaviors
From the pairwise correlation analysis results shown in Table III, when the APMV communicates with pedestrians using eHMI-T, there was only a significant correlation relationship was observed between the passengers’ Neuroticism scores and their gaze duration on the pedestrians’ head with a positive correlation. This result suggests that passengers with lower Neuroticism tended to avert their gaze from the pedestrian’s head while immersed in the interaction with the pedestrian and the silent eHMI-T, whereas passengers with higher Neuroticism tended to allocate longer visual attention to the pedestrian’s head.
However, studies [23, 24] suggested that people with a high Neuroticism score tended to look away when co-communicators looked them in the eyes during a conversation or other social situations. We considered that this difference arises because the studies [23, 24] focused on face-to-face interactions, where mutual eye contact was more likely to occur. In our experiment, the pedestrian primarily interacted with the eHMI and did not look at the passenger throughout the process. Moreover, since passengers could not directly access the communication content from the eHMI-T when the APMV interacted with pedestrians (study [1] also highlighted this issue via subjective evaluations), they had to infer it based on the pedestrians’ reactions. Given this uncertainty, we considered that passengers with higher neuroticism scores might remain more vigilant in such an ambiguous communication environment. Studies [25, 26] provided some thread showing that people with anxious personalities, such as high Neuroticism, exhibit increased vigilance toward potentially threatening stimuli.
@@ V-B Effects of eHMI-NV on passenger gaze behaviors
From the results of the pairwise correlation analysis shown in Table IV, when using eHMI-NV, no significant correlation was found between passengers’ Big Five personality traits and their gaze behavior. This result suggested that the neutral communication style of eHMI-NV provided a relatively consistent experience for passengers with different personality traits, and it was difficult to affect the gaze behavior of passengers with specific personality traits.
@@ V-D Limitations
The participants in this study are young Japanese, and the cultural and age factors may have influenced their specific gaze behavior. With only 24 participants, the small sample size may limit the statistical power of the pairwise correlation results. The results may be limited by the specific eHMI designs used in this study. The pedestrians in the experiment were played by experimenters. Their behaviors and reactions in the interaction scenarios were not as varied as those of real pedestrians. This limitation in diversity may potentially affect the results of this study.
@@ VI CONCLUSION
This study investigated the differences in gazing behavior towards pedestrians’ head by APMV passengers with different personalities when APMV communicates with pedestrians using three different types of eHMI When using a visual-based eHMI, which caused passengers to struggle in perceiving the communication content, the results indicated that passengers with higher Neuroticism scores, who were more sensitive to communication details, might seek cues from pedestrians’ reactions. In addition, eHMI-NV using neutral voice did not significantly affect the gaze behavior of passengers toward pedestrians, regardless of personality traits. In contrast, eHMI-AV using affective voice encouraged passengers with high Openness to Experience scores to focus on pedestrians’ heads.